# Optimal Parallel Algorithms for Convex Hulls in 2D and 3D under Noisy Primitive Operations

### v1

@@ E.1.2 Pruning halfspaces
Each halfspace $h\in H_{|C_{i}}^{nc}$ can determine whether it should be pruned in $O(\log n)$ time w.h.p. by projecting itself onto each face and performing a noisy binary searches on each contour to determine the closest hull point to the projected line representing $h$ (recall that we maintain the upper and lower hulls of each contour as a binary search tree, something we took advantage of to perform failure sweeping in Section 4). We can associate each contour point $p$ with the two halfspaces that define it so that they can be recovered in constant time. We can then perform a constant number of noisy comparisons using the trivial repetition strategy to determine which halfspace the ray passes through first in $O(\log n)$ time w.h.p. By the fact that our sample $R$ is good, $|H^{nc}|=O(n)$. Then we can perform each test in parallel in $O(\log n)$ span and $O(n\log n)$ work.
Now we consider the halfspaces that contribute to contours, which we denote as $H_{|C_{i}}^{c}$. Amato, Goodrich, and Ramos [5] show that, if a halfspace in $h\in H_{|C_{i}}^{c}$ contributes an edge to at least two of the contours, then it can be pruned. To determine which halfspaces should be pruned, we simply label each contour edge with the corresponding halfspace that generated it, which can be done when constructing the contours, and sort the labels lexicographically. Again, by our use of a good sample, the total number of contour halfspaces $|H^{c}|=O(n)$, thus we can perform noisy sorting in parallel over all $C_{i}$ in $O(\log n)$ span and $O(n\log n)$ work. In the lexicographically sorted array, if a halfspace is adjacent to itself, then it must contribute an edge to at least two contours, and so should be pruned. In [5], the authors show that after pruning, the set of all contour halfspaces across all cones and all subproblems also has size $O(n_{0})$.
@@ E.2 Failure Sweeping
We have shown that each step of the algorithm of [27], with the pruning method of [5], can be replaced with noise-tolerant versions that individually succeed with high probability in the current problem size. However, as we did for 2D convex hulls, we must negotiate the fact that our success probabilities get weaker as problem size gets smaller. To compensate for this, we once again perform failure sweeping. Unlike for 2D convex hulls, subproblems may not be evenly sized. We will denote $m$ as the size of the problem at recursive level $l$ and $M$ the size of its parent subproblem at recursive level $l-1$.
@@ E.3 Deferred Proofs
We will prove this by induction. Say we are at a parent subproblem of size $M$. Our inductive hypothesis assumes that we can compute each of its child subproblems of size $m$ with failure probability at most $m^{-c^{\prime}}$ in $O(\log m)$ span and $O(m\log m)$ work.
Our Section E.1 shows that the remaining steps to split and recombine subproblems take $O(\log M)$ span and $O(M\log M)$ work in total. From our discussion in the previous section, it takes $O(\log M)$ span and $O(m\log M)$ work to perform size reduction on larger subproblems. The total size of all subproblems of $M$ is at most $O(M)$ by pruning, therefore it takes $O(\log M)$ span and $O(M\log M)$ work in total to perform size reduction and the corresponding reconstruction after returning from the recursive calls w.h.p. in $M$. Lastly, Theorem 5.2, Lemma 5.3, Theorem 5.8, and our discussion of size reduction, it takes $O(\log M)$ span and $O(M\log M)$ work to failure sweep and recompute all failed subproblems w.h.p. in $M$. This holds as long as $c^{\prime}\geq 17$.
It remains to determine the failure parameter $c$ for each individual primitive operation. As noted at the end of Section E.1, sampling, constructing the sampled hull and point location structure, and pruning all take $O(M)$ noisy operations in total. In our discussion of size reduction, we observed that each level of reduction takes $O(m)$ noisy operations. Therefore, splitting and recombining problems takes $O(m)$ noisy operations in total, so $O(M)$ noisy operations over all of $M$’s subproblems. By Theorem 5.2 and Theorem 5.8, it takes $O(M)$ total noisy operations to failure sweep all problems and recompute any failed subproblems by brute-force. By union bound over all $O(M)$ invocations of noisy primitives, if a single noisy primitive fails with probability $\leq M^{-c}$, at least one fails with probability $O(M^{-(c-1)})$. Then we can take $c\geq c^{\prime}+1\geq 18$ to achieve the desired confidence bound.
@@
$T(n)=T(n^{9/10})+O(\log n),$
which evaluates to $O(\log n)$ total span as it follows a geometric sequence. Total work is proportional to problem size times span at each level of recursion. By pruning, each level of recursion has total problem size at most $O(n_{0})$, where $n_{0}$ is the initial problem size. Then work is bounded by the following recurrence:
$W(n)=W(n^{9/10})+O(n_{0}\log n)$
Again, this sums to $O(n_{0}\log n_{0})$. We conclude that it takes $O(\log n)$ span and $O(n\log n)$ work to compute all the intersection of $n$ halfspaces under noisy primitives.
Lemma E.22 (Same as Lemma 5.6).
Say that we have $O(M^{9/8})$ processes, each of which fail with high probability in $M^{1/8}$. Then at most a constant number of them fail with high probability in $M$.
Proof E.23.
We expect $O(M^{9/8})/M^{c^{\prime}/8}$ subproblems to fail, for some $c^{\prime}$ s.t. $c^{\prime}/8>1$. Let $X$ be a random variable representing the number of failed subproblems. Then, for any integer $q\geq 1$, Markov’s inequality states that $\Pr[X\geq q]\leq O(M^{9/8})/M^{c^{\prime}/8}\times 1/q\leq O(1/M^{(c^{\prime}-9)/8})$. If we set $c^{\prime}\geq 17$, this is a high-probability bound in $M$.

### v2

@@ E.1.2 Pruning halfspaces
Each halfspace $h\in H_{|C_{i}}^{nc}$ can determine whether it should be pruned in $O(\log n)$ time w.h.p. by projecting itself onto each face and performing a noisy binary searches on each contour to determine the closest hull point to the projected line representing $h$ (recall that we maintain the upper and lower hulls of each contour as a binary search tree, something we took advantage of to perform failure sweeping in Section 4). We can associate each contour point $p$ with the two halfspaces that define it so that they can be recovered in constant time. We can then perform a constant number of noisy comparisons using the trivial repetition strategy to determine which halfspace the ray passes through first in $O(\log n)$ time w.h.p. By the fact that our sample $R$ is good, $|H^{nc}|=O(n)$. Then we can perform each test in parallel in $O(\log n)$ span and $O(n\log n)$ work w.h.p.
Now we consider the halfspaces that contribute to contours, which we denote as $H_{|C_{i}}^{c}$. Amato, Goodrich, and Ramos [5] show that, if a halfspace in $h\in H_{|C_{i}}^{c}$ contributes an edge to at least two of the contours, then it can be pruned. To determine which halfspaces should be pruned, we simply label each contour edge with the corresponding halfspace that generated it, which can be done when constructing the contours, and sort the labels lexicographically. Again, by our use of a good sample, the total number of contour halfspaces $|H^{c}|=O(n)$, thus we can perform noisy sorting in parallel over all $C_{i}$ in $O(\log n)$ span and $O(n\log n)$ work w.h.p. In the lexicographically sorted array, if a halfspace is adjacent to itself, then it must contribute an edge to at least two contours, and so should be pruned. In [5], the authors show that after pruning, the set of all contour halfspaces across all cones and all subproblems also has size $O(n_{0})$.
@@ E.2 Failure Sweeping
We have shown that each step of the algorithm of [27], with the pruning method of [5], can be replaced with noise-tolerant versions that individually succeed with high probability in the current problem size. However, as we did for 2D convex hulls, we must negotiate the fact that our success probabilities get weaker as problem size gets smaller. To compensate for this, we once again perform failure sweeping. Unlike for 2D convex hulls, subproblems may not be evenly sized. We will denote $m$ as the size of the problem at recursive level $l$ and $M$ the size of its parent subproblem at recursive level l1l-1.
@@ E.3 Deferred Proofs
We will prove this by induction. Say we are at a parent subproblem of size $M$. Our inductive hypothesis assumes that we can compute each of its child subproblems of size $m$ with failure probability at most $M^{-c^{\prime}/2}$ in $O(\log M^{1/2})$ span and $O(m\log M^{1/2})$ work.
Our Section E.1 shows that the remaining steps to split and recombine subproblems take $O(\log M)$ span and $O(M\log M)$ work in total. From our discussion above, it takes $O(\log M)$ span and $O(M\log M)$ work to perform size reduction and the corresponding reconstruction after returning from the recursive calls w.h.p. in $M$. Lastly, by Theorem 5.2, Theorem 5.4, and our discussion of size reduction, it takes $O(\log M)$ span and $O(M\log M)$ work to failure sweep and recompute all failed subproblems w.h.p. in $M$. This holds as long as $c^{\prime}>21/8$.
It remains to determine the failure parameter $c$ for each individual primitive operation. As noted at the end of Section E.1, sampling, constructing the sampled hull and point location structure, and pruning all take $O(M)$ noisy operations in total. Our analysis above also implies that size reduction overall also takes $O(M)$ noisy operations. By Theorem 5.2 and Theorem 5.4, it takes $O(M)$ total noisy operations to failure sweep all problems and recompute any failed subproblems by brute-force. By union bound over all $O(M)$ invocations of noisy primitives, if a single noisy primitive fails with probability $\leq M^{-c}$, at least one fails with probability $O(M^{-(c-1)})$. Then we can take $c\geq c^{\prime}+1>29/8$ to achieve the desired confidence bound.
@@
$T(n)=T(\sqrt{n})+O(\log n),$
which evaluates to $O(\log n)$. Total work is proportional to problem size times span at each level of recursion. By pruning, each level of recursion has total problem size at most $O(n_{0})$, where $n_{0}$ is the initial problem size. Then work is bounded by the following recurrence:
$W(n)=W(\sqrt{n})+O(n_{0}\log n).$
This sums to $O(n_{0}\log n_{0})$. We conclude that it takes $O(\log n)$ span and $O(n\log n)$ work w.h.p. to compute all the intersection of $n$ halfspaces under noisy primitives.